# Cap2Sum: Learning to Summarize Videos by Generating Captions

### v1

@@ 1 Introduction
Due to the rapid expansion of video data on the Internet today, video summarization algorithm becomes very useful to users. For example, for those videos recorded by wearable devices that are hours long, it is meaningful to extract the key frames in them. However, the cost of labelling video summarization datasets is extremely high. Specifically, the annotators have to label the importance score for each frame in the videos, and each video requires multiple annotators to reduce subjectivity. Therefore, the existing video summarization datasets (e.g., TVSum Song et al. (2015), SumMe Gygli et al. (2014)) are all small-scale datasets, which are insufficient to support the research of video summarization based on deep learning. This fact leads to the limited performance and generalization capacity of existing supervised video summarization methods. Park et al. (2020); Gygli et al. (2015); Zhang et al. (2016b); Rochan et al. (2018); Zhang et al. (2016a); Zhao et al. (2017); Zhang et al. (2018).
@@
To address the insufficiency of large-scale training data, unsupervised and weakly-supervised video summarization methods Yuan et al. (2019); Mahasseni et al. (2017); Zhou et al. (2018a); Ho et al. (2018); Panda et al. (2017); Cai et al. (2018) have been investigated recently. Generally speaking, previous unsupervised approaches can be divided into (1) adversarial-learning-based methods Apostolidis et al. (2020); Jung et al. (2020, 2019) and (2) reinforcement-learning-based methods Zhou et al. (2018a). For (1), the video summarization model (i.e., summarizer) is trained to generate a summarization that can reconstruct the original input video. For (2), the summarization is performed by targeting some specific properties (e.g., diversity, length) of an optimal summary. However, for (1), the reconstruction of the original video requires less important information (e.g., backgrounds, environment) contains in the summary, thus is not an ideal goal for training a video summarizer; for (2), the optimization targets are very heuristic. Instead of completely abandoning the usage of any ground-truth labels, weakly-supervised based methods exploit less-expensive weak labels (e.g., video categories, video metadata) to train the summarizers. Unfortunately, due to the insufficiency of the weak labels used, these weakly-supervised methods do not have performance advantages compared to the supervised-based methods. But we still see potential in weakly-supervised methods, as these methods can be trained on large-scale datasets under reasonable supervision.
@@
We notice that a key factor of the weakly-supervised video summarization is the choice of weak labels. A reasonable label should (a) be easy to labelled, and (b) contain extensive information on video summarization. Based on the above requirements, we ask if we can generate video captions as weak labels to train a strong video summarizer. For (a), video captions are easy to be annotated and therefore many large-scale datasets Xu et al. (2016); Zhou et al. (2018b); Krishna et al. (2017) are currently available. For (b), video captions can also be regarded as the summarization of video contents, but in text form rather than as frame-level scores, making it ideal for training video summarizers. Inspired by this, in this paper, we introduce Cap2Sum, to train video summarizers effectively via adopting video captions as weak labels. Specifically, the proposed model contains two main transformer-based components, a video summarizer, and a video captioner. During training, given a video with video caption annotations, the video summarizer summarizes the video and generates frame-level scores. We use these scores to weigh the video features and feed them to the video captioner to generate dense captions (i.e., text captions with start time and end time). The model is trained and supervised by ground-truth captions. To generate accurate captions, the summarizer is expected to find the frames that are related to the captions, which actually contain the key content of the video, approximating a video summary. In practical, Cap2Sum can generate a summary in a zero-shot setting or be fine-tuned by the labels in the target dataset. The labels employed for fine tuning can either be the ground-truth summary or the video caption. Therefore, to examine the performance of Cap2Sum under the weakly-supervised fine-tuning using video caption, we proposed the TVSum-Caption and SumMe-Caption dataset based on the commonly used video summarization dataset TVSum Song et al. (2015) and SumMe Gygli et al. (2014). The proposed datasets will be publicly released for further research.
@@
In practice, due to the huge domain gap between videos from the video caption datasets and the video summarization datasets (videos from video caption datasets are mostly about the activities of characters and are mostly in third-person view, while videos from video summarization datasets may not contain characters and include first-person videos), the zero-shot performance is not promising for the models pertained on video caption datasets. To address this problem, we propose a CLIP prior mechanism to complement the learning of out-of-distribution objects. This mechanism exploits CLIP Radford et al. (2021) to find the frames which contain non-character objects with high information entropy in the videos. We think these frames should be included in the summary and generating auxiliary supervision for training. We conduct extensive experiments on TVSum, SumMe, and the proposed datasets. The ablation studies demonstrate the effectiveness of the proposed Cap2Sum architecture and CLIP-Prior mechanism. Our approach achieves significant performance improvement compared to the relevant method CLIP-It Narasimhan et al. (2021) (4.5% F1-Score improvement on SumMe using Transfer settings), and reaches state-of-the-art performance among current supervised and unsupervised methods.
@@ 2.1 Video Summarization
The goal of video summarization is to create a condensed version of the source video that conveys sufficient information and highlights significant events. Current methods can be categorized into three groups: supervised, unsupervised, and weakly-supervised learning approaches. Supervised learning approaches Zhang et al. (2016a, b); Zhou et al. (2018a); Zhang et al. (2018); Fajtl et al. (2018); Park et al. (2020); Zhao et al. (2021); Narasimhan et al. (2021); Zhu et al. (2022); Jiang and Mu (2022); He et al. (2023) learn from raw videos with corresponding ground truth summary videos annotated by humans. With the help of the benchmark datasets (e.g., SumMe Gygli et al. (2014) and TVSum Song et al. (2015)), supervised learning approaches have achieved impressive performance. Among them, CLIP-It Narasimhan et al. (2021) utilizes an off-the-shelf video captioning model and CLIP Radford et al. (2021) (a large-scale vision-language model) to generate summary videos conditioned on the text. iPTNet Jiang and Mu (2022) employs a joint training approach for both the video summarization and correlated moment localization tasks, which leverages extra moment localization data samples to enhance the performance of video summarization. A2Summ He et al. (2023) builds a unified transformer-based framework for multimodal summarization. A significant drawback of supervised video summarization is its reliance on labeled training data. Typically, the datasets used in this field are created by human annotators to access the input video and identify the key frames or shots. Therefore, this highly resource-intensive annotation process leads to the lack of large-scale benchmark datasets for video summarization.
Unsupervised learning approaches Mahasseni et al. (2017); Li et al. (2019); Rochan et al. (2018); Rochan and Wang (2019); He et al. (2019); Park et al. (2020); Wu et al. (2021); Phaphuangwittayakul et al. (2021); Apostolidis et al. (2022) utilize various hand-crafted heuristics to measure the score of video frames. For example, Mahasseni et al. Mahasseni et al. (2017) propose a generative architecture based on variational recurrent auto-encoders and generative adversarial networks to select a subset of key frames. ACGAN He et al. (2019) utilizes attentive conditional GANs, with a generator that produces weighted frame features and importance scores, and a discriminator that distinguishes between weighted and raw frame features. Weakly-supervised learning approaches Ho et al. (2018); Panda et al. (2017); Cai et al. (2018) in video summarization often exploit auxiliary video information such as video categories or metadata. Although unsupervised and weakly-supervised learning methods can alleviate the problem of insufficient summary annotations to some extent, it is difficult to design flawless heuristics for the former, while the latter still rely on auxiliary video information, making their performance hard to match that of supervised learning methods.
@@ 2.2 Discussion
When we train a video summarizer, we actually expect our model to learn the human prior knowledge of what humans consider “important" in a video. For the supervised learning, this prior knowledge of “importance" contains in the ground-truth label (i.e., the importance score under human annotation) and is easy to learn. But for unsupervised or weakly-supervised learning, we have to find other ways to let the model learn this human prior. The promising performance improvement achieved by CLIP-It Narasimhan et al. (2021) has demonstrated that the video captions do relate to the video summarization. However, the improvement in CLIP-It is limited by the accuracy of the generated captions. According to our tests, the caption model BMT Iashin and Rahtu (2020) used in CLIP-It performs poorly on the video summarization datasets (i.e., TVSum and SumMe). Besides, the training of CLIP-It is still conducted on the small-scale dataset, leading to the limited zero-shot performance. Generally speaking, compared to CLIP-It, the proposed Cap2Sum has following improvements and novelties: (1) Instead a combining of all the captions in a long video, we propose to use the dense video captions, which are more accurate for the learning of important events in the videos. (2) CLIP-It uses the video captions generated by off-the-shelf models as extra inputs while we adopt the video captions as weak labels, that means the proposed Cap2Sum can be pre-trained on large-scale dense video caption datasets and tested on any videos without any annotation. (3) In our method, the video summarizer can generate captions without video captioner during testing. Thus, the performance of Cap2Sum is not limited by the off-the-shelf pre-trained video captioners.
@@ 3 Methodology
As we mentioned in Section 1, the extremely high labelling cost of video summarization motivates us to find other weak labels to train the summarizers. In this section, we propose Cap2Sum, a weakly-supervised video summarization model trained by dense video captions. The major component of Cap2Sum is a video summarizer and a video captioner. Cap2Sum is trained on the dense video caption datasets and expects the model to learn which frames are related to the caption of a video. Through the weakly-supervised training on large-scale datasets, our model achieves strong performance and generalization capacity.
@@ 3.1 Overview of Cap2Sum
In this section, we introduce the architecture of Cap2Sum and how model is trained under the supervision of dense video captions. As shown in Fig. 1 (a), the proposed Cap2Sum is composed of a visual encoder $f_{enc}$, a video summarizer $f_{sum}$, a pre-trained CLIP $f_{clip}$, and a pre-trained video captioner $f_{cap}$. During training, for a given video $\mathbf{V}\in\mathbb{R}^{T\times C\times H\times W}$ with dense caption annotation $c=\{(s_{i},e_{i},t_{i})\}$, the video frames are encoded by an image encoder (we employ the image encoder of CLIP Radford et al. (2021) in this paper) to generate frame features $\mathbf{F}=f_{enc}(\mathbf{V})$. A transformer-based video summarizer generates the summary (i.e., frame-wise scores) $\mathbf{S}\in\mathbb{R}^{T}$ based on the visual features. These scores are used to produce weighted features $\mathbf{F}_{w}=\mathbf{S}\cdot\mathbf{F}$, where $\cdot$ denotes element-wise multiplication. Then, the weighted feature is fed to the video captioner to predict captions. As an auxiliary supervision, we propose a CLIP Prior Generator (CLIP Prior Gen.) to generate the CLIP Prior. We introduce this mechanism in detail in Section 3.2. During testing, for the input videos without annotation, Cap2Sum can perform zero-shot video summarization. To improve the performance, Cap2Sum can also perform few-shot video summarization via the supervised fine-tuning or weakly-supervised fine-tuning on target datasets. Following previous works Narasimhan et al. (2021); Apostolidis et al. (2020); Jung et al. (2020), we use 0/1 knapsack algorithm Zhao and Xing (2014) to select the frames that should be included in the final summary. The details of Cap2Sum are as follows.
@@
Video Summarizer. The key component to perform the summarization is the video summarizer. To generate a reasonable summary, the summarizer has to understand the video contents and combine the context information. Based on the above requirements, we decided to use the video transformer, which has a good capacity to capture global temporal relationships, as the summarizer. As shown in Fig. 1 (b), the frame features are first embedded to the token vectors, and then fed to the transformer blocks. The transformer blocks conduct multi-head self-attention (MHSA) to the tokens. Finally, a MLP head transfer the token sequence to the frame-wise scores.
@@
Video Captioner. The captioner employed in Cap2Sum is used to support the training of the summarizer. To obtain the best gradients, we prefer to use a high-performance dense video caption model. In this paper, we apply PDVC Wang et al. (2021) as the captioner of Cap2Sum. PDVC is an end-to-end dense video captioner based on deformable transformer. As shown in Fig. 1 (c), PDVC captioner is composed of a transformer encoder-decoder network, an event counter, a localization head and a caption head. The frame features are encoded by the encoder, and then decoded conditioned by the decoder. The event counter predicts the event count in the video, the localization and caption head generates the captions with segments. In practice, the choice of captioner has little impact on the performance of Cap2Sum, and Cap2Sum is able to achieve promising performance on various captioner with acceptable performance.
@@ 3.2 CLIP Prior Mechanism
In practice, the videos in the video caption datasets (most are third-person and contain characters) have huge domain gap with the videos in the video summarization datasets (most are first-person and may not contain characters). Besides, the ground-truth captions tend to focus on the activities of the characters and ignore important objects that are not related to the activities. Therefore, when testing on the video summarization datasets under zero-shot setting, the model is hard to correctly label the frames which contain important objects (i.e., fire on the food, bus in the tunnel) as key frames. To address this problem, we propose a CLIP Prior mechanism that exploits the prior knowledge learnt by CLIP Radford et al. (2021). Specifically, CLIP Prior is generated by the CLIP Prior Generator shown in Fig. 2. We select 100 common object labels mainly based on the categories of CIFAR-100 dataset Krizhevsky et al. (2009). We then use CLIP to check whether each object exists in each frame, by calculating the similarity between frames and text. Specifically, we use the prompt “An image of [object].", where [object] denote the object label, to construct 100 sentences. These pre-defined texts are embedded by the text encoder of CLIP, and then calculate the similarity with the features obtained by the CLIP image encoder (i.e., frame features in Fig. 1) for each frame via matrix multiplication, formulated as:
@@ A. Weakly-supervised pre-training
Caption Loss. The major loss function to exploit the caption annotation is the caption loss. Following Wang et al. (2021), the caption loss is composed of $L_{giou}$ (i.e., the generalized IOU Rezatofighi et al. (2019) between predicted temporal segments and ground-truth segments), $L_{cls}$ (i.e., focal loss Lin et al. (2017) between the prediction of prediction scores and the ground-truth), $L_{ec}$ (i.e., the predicted event count and the ground-truth), and $L_{pred}$ (i.e., the cross-entropy between the caption prediction and the ground-truth caption):
@@
CLIP Prior Loss. As mentioned in Section 3.2, we introduce the CLIP Prior mechanism to enhance the generalization capacity of Cap2Sum. Specifically, we employ a mean-squared error (MSE) loss between the CLIP Prior and masked summarization, formulated as:
@@
However, our method can be fine-tuned by the captions of the videos in the testing dataset. In this scenario, the fine-tuning is weakly-supervised by the caption loss same to Eq. 2.

### v2

@@ 1 Introduction
Due to the rapid expansion of video data on the Internet today, video summarization algorithm becomes very useful to users. For example, for those videos recorded by wearable devices that are hours long, it is meaningful to extract the key frames in them. However, the cost of labelling video summarization datasets is extremely high. Specifically, the annotators have to label the importance score for each frame in the videos, and each video requires multiple annotators to reduce subjectivity. Therefore, the existing video summarization datasets (e.g., TVSum [29], SumMe [5]) are all small-scale datasets, which are insufficient to support the research of video summarization based on deep learning. This fact leads to the limited performance and generalization capacity of existing supervised video summarization methods [23, 6, 41, 28, 40, 42].
@@
To address the insufficiency of large-scale training data, unsupervised and weakly-supervised methods [39, 20, 45, 9, 22, 33, 32, 3] have been investigated recently. Generally speaking, previous unsupervised approaches can be divided into (1) adversarial-learning-based methods [1, 13, 12, 30, 34] and (2) reinforcement-learning-based methods [45]. For (1), the video summarization model (i.e., summarizer) is trained to generate a summarization that can reconstruct the original input video. For (2), the summarization is performed by targeting some specific properties (e.g., diversity, length) of an optimal summary. However, for (1), the reconstruction of the original video requires less important information (e.g., backgrounds, environment) contains in the summary, thus is not an ideal goal for training a video summarizer; for (2), the optimization targets are very heuristic. Instead of completely abandoning the usage of any ground-truth labels, weakly-supervised based methods exploit less-expensive weak labels (e.g., video categories, video metadata) to train the summarizers. Unfortunately, due to the insufficiency of the weak labels used, these weakly-supervised methods do not have performance advantages compared to the supervised-based methods. But we still see potential in weakly-supervised methods, as these methods can be trained on large-scale datasets under reasonable supervision.
@@
We notice that a key factor of the weakly-supervised video summarization is the choice of weak labels. A reasonable label should (a) be easy to labelled, and (b) contain extensive information on video summarization. Based on the above requirements, we ask if we can generate video captions as weak labels to train a strong video summarizer. For (a), video captions are easy to be annotated and therefore many large-scale datasets [38, 46, 16] are currently available. For (b), video captions can also be regarded as the summarization of video contents, but in text form rather than as frame-level scores, making it ideal for training video summarizers. Inspired by this, in this paper, we introduce Cap2Sum, to train video summarizers effectively via adopting video captions as weak labels. Specifically, the proposed model contains two main transformer-based components, a video summarizer, and a video captioner. During training, given a video with video caption annotations, the video summarizer summarizes the video and generates frame-level scores. We use these scores to weigh the video features and feed them to the video captioner to generate dense captions (i.e., text captions with start time and end time). The model is trained and supervised by ground-truth captions. To generate accurate captions, the summarizer is expected to find the frames that are related to the captions, which actually contain the key content of the video, approximating a video summary. In practical, Cap2Sum can generate a summary in a zero-shot setting or be fine-tuned by the labels in the target dataset. The labels employed for fine tuning can either be the ground-truth summary or the video caption. Therefore, to examine the performance of Cap2Sum under the weakly-supervised fine-tuning using video caption, we proposed the TVSum-Caption and SumMe-Caption dataset based on the commonly used video summarization dataset TVSum [29] and SumMe [5]. The proposed datasets will be publicly released for further research.
@@
In practice, due to the huge domain gap between videos from the video caption datasets and the video summarization datasets (videos from video caption datasets are mostly about the activities of characters and are mostly in third-person view, while videos from video summarization datasets may not contain characters and include first-person videos), the zero-shot performance is not promising for the models pertained on video caption datasets. To address this problem, we propose a CLIP prior mechanism to complement the learning of out-of-distribution objects. This mechanism exploits CLIP [25] to find the frames which contain non-character objects with high information entropy in the videos. We think these frames should be included in the summary and generating auxiliary supervision for training. We conduct extensive experiments on TVSum, SumMe, and the proposed datasets. The ablation studies demonstrate the effectiveness of the proposed Cap2Sum architecture and CLIP-Prior mechanism. Our approach achieves significant performance improvement compared to the relevant method CLIP-It [21] (4.5% F1-Score improvement on SumMe using Transfer settings), and reaches state-of-the-art performance among current supervised and unsupervised methods.
@@ 2.1 Video Summarization
The goal of video summarization is to create a condensed version of the source video that conveys sufficient information and highlights significant events. Current methods can be categorized into three groups: supervised, unsupervised, and weakly-supervised learning approaches. Supervised learning approaches [40, 41, 45, 42, 4, 23, 44, 21, 47, 11, 7] learn from raw videos with corresponding ground truth summary videos annotated by humans. With the help of the benchmark datasets (e.g., SumMe [5] and TVSum [29]), supervised learning approaches have achieved impressive performance. Among them, CLIP-It [21] utilizes an off-the-shelf video captioning model and CLIP [25] (a large-scale vision-language model) to generate summary videos conditioned on the text. iPTNet [11] employs a joint training approach for both the video summarization and correlated moment localization tasks, which leverages extra moment localization data samples to enhance the performance of video summarization. A2Summ [7] builds a unified transformer-based framework for multimodal summarization. A significant drawback of supervised video summarization is its reliance on labeled training data. Typically, the datasets used in this field are created by human annotators to access the input video and identify the key frames or shots. Therefore, this highly resource-intensive annotation process leads to the lack of large-scale benchmark datasets for video summarization.
Unsupervised learning approaches [20, 18, 28, 31, 27, 8, 23, 37, 24, 2] utilize various hand-crafted heuristics to measure the score of video frames. For example, Mahasseni et al. [20] propose a generative architecture based on variational recurrent auto-encoders and generative adversarial networks to select a subset of key frames. ACGAN [8] utilizes attentive conditional GANs, with a generator that produces weighted frame features and importance scores, and a discriminator that distinguishes between weighted and raw frame features. Weakly-supervised learning approaches [9, 22, 3] in video summarization often exploit auxiliary video information such as video categories or metadata. Although unsupervised and weakly-supervised learning methods can alleviate the problem of insufficient summary annotations to some extent, it is difficult to design flawless heuristics for the former, while the latter still rely on auxiliary video information, making their performance hard to match that of supervised learning methods.
@@ 2.2 Discussion
When we train a video summarizer, we actually expect our model to learn the human prior knowledge of what humans consider “important" in a video. For the supervised learning, this prior knowledge of “importance" contains in the ground-truth label (i.e., the importance score under human annotation) and is easy to learn. But for unsupervised or weakly-supervised learning, we have to find other ways to let the model learn this human prior. The promising performance improvement achieved by CLIP-It [21] has demonstrated that the video captions do relate to the video summarization. However, the improvement in CLIP-It is limited by the accuracy of the generated captions. According to our tests, the caption model BMT [10] used in CLIP-It performs poorly on the video summarization datasets (i.e., TVSum and SumMe). Besides, the training of CLIP-It is still conducted on the small-scale dataset, leading to the limited zero-shot performance. Generally speaking, compared to CLIP-It, the proposed Cap2Sum has following improvements and novelties: (1) Instead a combining of all the captions in a long video, we propose to use the dense video captions, which are more accurate for the learning of important events in the videos. (2) CLIP-It uses the video captions generated by off-the-shelf models as extra inputs while we adopt the video captions as weak labels, that means the proposed Cap2Sum can be pre-trained on large-scale dense video caption datasets and tested on any videos without any annotation. (3) In our method, the video summarizer can generate captions without video captioner during testing. Thus, the performance of Cap2Sum is not limited by the off-the-shelf pre-trained video captioners.
@@ 3 Methodology
As we mentioned in Section˜1, the extremely high labelling cost of video summarization motivates us to find other weak labels to train the summarizers. In this section, we propose Cap2Sum, a weakly-supervised video summarization model trained by dense video captions. The major component of Cap2Sum is a video summarizer and a video captioner. Cap2Sum is trained on the dense video caption datasets and expects the model to learn which frames are related to the caption of a video. Through the weakly-supervised training on large-scale datasets, our model achieves strong performance and generalization capacity.
@@ 3.1 Overview of Cap2Sum
In this section, we introduce the architecture of Cap2Sum and how model is trained under the supervision of dense video captions. As shown in Fig.˜1 (a), the proposed Cap2Sum is composed of a visual encoder $f_{enc}$, a video summarizer $f_{sum}$, a pre-trained CLIP $f_{clip}$, and a pre-trained video captioner $f_{cap}$. During training, for a given video $\mathbf{V}\in\mathbb{R}^{T\times C\times H\times W}$ with dense caption annotation $c=\{(s_{i},e_{i},t_{i})\}$, the video frames are encoded by an image encoder (we employ the image encoder of CLIP [25] in this paper) to generate frame features $\mathbf{F}=f_{enc}(\mathbf{V})$. A transformer-based video summarizer generates the summary (i.e., frame-wise scores) $\mathbf{S}\in\mathbb{R}^{T}$ based on the visual features. These scores are used to produce weighted features $\mathbf{F}_{w}=\mathbf{S}\cdot\mathbf{F}$, where $\cdot$ denotes element-wise multiplication. Then, the weighted feature is fed to the video captioner to predict captions. As an auxiliary supervision, we propose a CLIP Prior Generator (CLIP Prior Gen.) to generate the CLIP Prior. We introduce this mechanism in detail in Section˜3.2. During testing, for the input videos without annotation, Cap2Sum can perform zero-shot video summarization. To improve the performance, Cap2Sum can also perform few-shot video summarization via the supervised fine-tuning or weakly-supervised fine-tuning on target datasets. Following previous works [21, 1, 13], we use 0/1 knapsack algorithm [43] to select the frames that should be included in the final summary. The details of Cap2Sum are as follows.
@@
Video Summarizer. The key component to perform the summarization is the video summarizer. To generate a reasonable summary, the summarizer has to understand the video contents and combine the context information. Based on the above requirements, we decided to use the video transformer, which has a good capacity to capture global temporal relationships, as the summarizer. As shown in Fig.˜1 (b), the frame features are first embedded to the token vectors, and then fed to the transformer blocks. The transformer blocks conduct multi-head self-attention (MHSA) to the tokens. Finally, a MLP head transfer the token sequence to the frame-wise scores.
@@
Video Captioner. The captioner employed in Cap2Sum is used to support the training of the summarizer. To obtain the best gradients, we prefer to use a high-performance dense video caption model. In this paper, we apply PDVC [36] as the captioner of Cap2Sum. PDVC is an end-to-end dense video captioner based on deformable transformer. As shown in Fig.˜1 (c), PDVC captioner is composed of a transformer encoder-decoder network, an event counter, a localization head and a caption head. The frame features are encoded by the encoder, and then decoded conditioned by the decoder. The event counter predicts the event count in the video, the localization and caption head generates the captions with segments. In practice, the choice of captioner has little impact on the performance of Cap2Sum, and Cap2Sum is able to achieve promising performance on various captioner with acceptable performance.
@@ 3.2 CLIP Prior Mechanism
In practice, the videos in the video caption datasets (most are third-person and contain characters) have huge domain gap with the videos in the video summarization datasets (most are first-person and may not contain characters). Besides, the ground-truth captions tend to focus on the activities of the characters and ignore important objects that are not related to the activities. Therefore, when testing on the video summarization datasets under zero-shot setting, the model is hard to correctly label the frames which contain important objects (i.e., fire on the food, bus in the tunnel) as key frames. To address this problem, we propose a CLIP Prior mechanism that exploits the prior knowledge learnt by CLIP [25]. Specifically, CLIP Prior is generated by the CLIP Prior Generator shown in Fig.˜2. We select 100 common object labels mainly based on the categories of CIFAR-100 dataset [17]. We then use CLIP to check whether each object exists in each frame, by calculating the similarity between frames and text. Specifically, we use the prompt “An image of [object].", where [object] denote the object label, to construct 100 sentences. These pre-defined texts are embedded by the text encoder of CLIP, and then calculate the similarity with the features obtained by the CLIP image encoder (i.e., frame features in Fig.˜1) for each frame via matrix multiplication, formulated as:
@@ A. Weakly-supervised pre-training
Caption Loss. The major loss function to exploit the caption annotation is the caption loss. Following [36], the caption loss is composed of $L_{giou}$ (i.e., the generalized IOU [26] between predicted temporal segments and ground-truth segments), $L_{cls}$ (i.e., focal loss [19] between the prediction of prediction scores and the ground-truth), $L_{ec}$ (i.e., the predicted event count and the ground-truth), and $L_{pred}$ (i.e., the cross-entropy between the caption prediction and the ground-truth caption):
@@
CLIP Prior Loss. As mentioned in Section˜3.2, we introduce the CLIP Prior mechanism to enhance the generalization capacity of Cap2Sum. Specifically, we employ a mean-squared error (MSE) loss between the CLIP Prior and masked summarization, formulated as:
@@
However, our method can be fine-tuned by the captions of the videos in the testing dataset. In this scenario, the fine-tuning is weakly-supervised by the caption loss same to Eq.˜2.